c.) Suppose a monopoly producer is also a monopsonist in the labour market. Demand for the output is P=100 - Q. The production function is Q= L, and the labour supply curve is w= 10 + L. How much labour does the firm hire? What wage is paid? [10 marks]




d.) Suppose the labour market is competitive, the supply curve of labour is upward sloping and the amount of capital is fixed. If the output market changes from a competitive market to a monopoly, what is the effect on its demand for labour? Explain briefly (in not more than 50 words). [7 marks]

1
Expert's answer
2021-11-17T10:00:40-0500

(c)

"MRP=100-2L"

Marginal Expenditure"=10+2L"

Marginal Expenditure"=MRP"

"10+2L=100-2L"

"L=\frac{90}{4}"

"=22.5" units of labor.

"w=10+L"

"w=10+22.5=32.5"


(d)

In a competitive labor market, the equilibrium wage and employment level are determined at the point where the demand for labor equals the supply of labor. This is as shown below:



Under the influence of monopoly power, prices paid by consumers will raise and national income will shift away from workers and to corporate profits. In case the labor market concentration affects all workers equally, there won't be increase in wage or compensation inequality.